OpenStudy (anonymous):

The measures of each exterior angle in a regular hexagon is (4x + 15)°. Find the value of x. Show equations and all work that leads to your answer.

OpenStudy (anonymous):

exterior angle in regular hexagon is 360 /6 = 60 degrees so 4x + 15 = 60

OpenStudy (anonymous):

4x = 60 - 15 = 45 x = 45/4 = 11.25 degrees